OpenStudy (anonymous):

HELP PLEASE LAST MINUTE WORK NEED NOW> The equation of line EF is y = 1 over 2x + 6. Write an equation of a line parallel to line EF in slope-intercept form that contains point (0, −2). y = −2x − 2 y = negative 1 over 2x + 2 y = 1 over 2x − 2 y = − 2x + 2

OpenStudy (anonymous):

@pooja195 @e.mccormick @Nnesha

OpenStudy (anonymous):

y = 2x + 2

Parth (parthkohli):

that does not contain the point (0,-2), heartless

OpenStudy (anonymous):

No, Im very comfused..

OpenStudy (e.mccormick):

@II-Heartless-ll 1) That is not one of the answers. 2) Do not do answers only... read the CoC. http://openstudy.com/code-of-conduct

OpenStudy (e.mccormick):

For the other line to be parallel it must have the same slope. So all you need is a line with the same slope that has that point.

OpenStudy (anonymous):

I understand the COC I need help though, And I have very limited time right now.

OpenStudy (e.mccormick):

I was not taling to you, Lex, about the CoC. =)

OpenStudy (texaschic101):

what answer choice has the same slope as the original equation ?

OpenStudy (anonymous):

I would guess either y=2x+2 or 1 over 2x-2 ... im not very good at this subject though :/

Nnesha (nnesha):

y = mx +B slope intercept form m = slope b =y-intercept

OpenStudy (anonymous):

Would it be A then.

OpenStudy (texaschic101):

there is only 1 answer choice that has the same slope as the original equation

OpenStudy (e.mccormick):

I am betting your y = 1 over 2x + 6 means: \(y = \dfrac{1}{2}x + 6\) So you need the same 1/2 slope and then just put in the points with b to find the new intercept,.

OpenStudy (e.mccormick):

Oooh! Good point Tex! Yah, that is a huge givaway!

Parth (parthkohli):

wait, wasn't that obvious

OpenStudy (texaschic101):

not necessary...only 1 answer choice has the same slope....parallel lines have same slope...the other 3 answers are automatically sliminated

OpenStudy (texaschic101):

typo...eliminated

OpenStudy (e.mccormick):

Parth, yah... with it being the only one with the same slope, the entire question is a but of a giveaway.

OpenStudy (anonymous):

This is confusing for me, im sorry, would it be B

OpenStudy (e.mccormick):

a bit...

OpenStudy (texaschic101):

what is the slope of the original equation

OpenStudy (e.mccormick):

Lex, you slowp is 1/2 or 1 over 2. Which answer has that slope?

OpenStudy (texaschic101):

lol....and you said I gave a big hint

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha y = mx +B slope intercept form m = slope b =y-intercept \(\color{blue}{\text{End of Quote}}\) m is slope which is ALWAYS with x o^_^o ;p

OpenStudy (anonymous):

C!

OpenStudy (anonymous):

Thank you guys sorry im a slow poke.

OpenStudy (texaschic101):

BINGO :)

OpenStudy (e.mccormick):

It is OK. In the end you saw the key part.

OpenStudy (anonymous):

Ive been enlightened by this :3

OpenStudy (e.mccormick):

Good Grace.

OpenStudy (anonymous):

:3